OpenStudy (anonymous):

The emf of the cell Ag | AgCl (saturated solution) || Cl-(c1 M)|AgCl|Ag is given by: where Ksp is the solubility product of AgCl

OpenStudy (anonymous):

This is what I got but the answer in the book is -ve of this that is -E cell.

OpenStudy (vincent-lyon.fr):

Hi! It depends on convention when you write the symbolic cell. In France, E is E1 - E2 where side 1 is on the right and side 2 is on the left. With this convention, I find: E = 0.059 lg (√Ks/c1)

OpenStudy (anonymous):

Can you please check this? I took cathode(right one) concentration = that of Cl ions and anodic as SQRT(Ksp) Is that correct?

OpenStudy (anonymous):

Also, I have another question, how to determine the temperature dependence of EMF?

OpenStudy (anonymous):

I mean, you can take the second term as +ve or -ve based on the order of writing concentrations in the second term. So, how does temperature affect the EMF?

OpenStudy (vincent-lyon.fr):

I think it is definitely a matter of convention. In France we write Nernst equation with a + sign and ox activity in the numerator. E = E° + 0.059/n lg (a(ox)/a(red)) Then E1 = E°(Ag+/Ag) + 0.059 lg [Ag+]1 E2 = E°(Ag+/Ag) + 0.059 lg [Ag+]2 E = E1 - E2 = 0.059 lg ([Ag+]1/[Ag+]1) = 0.059 + lg \(\large \frac {K_s/c_1}{\sqrt K_s}\) = 0.059 + lg \(\large \frac {\sqrt K_s}{c_1}\)

OpenStudy (vincent-lyon.fr):

Remember where the 0.059 V comes from. At T = 298 K, \(\Large \frac {RT\ln10}{F}\) = 0.059V Simply change 298 for 308 K